OpenStudy (kpop4life123):

Which equation best represents the line? (4 points) y = 3/4 x + 4 y = 4/3 x + 4 y = 4x + 3/4 y = 4x + 4/3

OpenStudy (sparrow2):

1 and 2 are identical

OpenStudy (sparrow2):

it;s y=x+4

OpenStudy (kpop4life123):

y = 3/4 x + 4 y = 4/3 x + 4 y = 4x + 3/4 y = 4x + 4/3

OpenStudy (kpop4life123):

@sparrow2 ??